geerky42 (geerky42):

sec²x + csc²x = A) 2secxcscx B) csc2x C) 2secx D) 2csc(2x) E) sec²x·csc²x

OpenStudy (anonymous):

substitute for sec nd csc in terms of sin nd cos

geerky42 (geerky42):

\(\Large \displaystyle \frac{1}{\sin^2 x}+ \frac{1}{\cos^2 x}\) ··· Now what?

OpenStudy (anonymous):

take lcm and add them

geerky42 (geerky42):

\[\Huge \frac{\cos^2 x + \sin^2 x}{\sin^2 x \cos^2 x}\]

OpenStudy (anonymous):

cos^2x+sin^2x = ?

geerky42 (geerky42):

E.